OpenStudy (samigupta8):

The condition that the parabolas y^2=4c(x-d) and y^2=4ax have a common normal other than x axis (a>0,c>0 and c>a) is

OpenStudy (samigupta8):

@ganeshie8

OpenStudy (samigupta8):

I found this a/m=-md+c/m

OpenStudy (samigupta8):

@phi

OpenStudy (phi):

what's a "common normal" ?

OpenStudy (samigupta8):

The normal that is common to both the curves!

OpenStudy (samigupta8):

Options are A).2a<2c+d B).2c<2a+d C).2d<2a+c D).2d<2c+a

OpenStudy (phi):

yes, but what is the definition of a normal ?

OpenStudy (samigupta8):

A line that is perpendicular to tangent.

OpenStudy (samigupta8):

@ikram002p

OpenStudy (samigupta8):

@pawanyadav

OpenStudy (pawanyadav):

Suppose a point (x,y) Find dy/dx for both curve then equate it..

OpenStudy (samigupta8):

Did u see my equation in the second post ??

OpenStudy (pawanyadav):

What is m there ?

OpenStudy (pawanyadav):

slope of tangent or normal

OpenStudy (samigupta8):

Wait!! I did wrong.. I equated the equations of common tangent to the curves

OpenStudy (samigupta8):

I got this now 2am+am^3=2cm+cm^3+md

OpenStudy (pawanyadav):

So you got option A

OpenStudy (samigupta8):

Oops A only!

OpenStudy (samigupta8):

Ty

OpenStudy (pawanyadav):

Yes. ..

OpenStudy (samigupta8):

No ...it should be 2a>2c+d

OpenStudy (samigupta8):

@pawanyadav

OpenStudy (pawanyadav):

Why..

OpenStudy (samigupta8):

I got this 2a+am^2=2c+cm^2+d And if we take the terms to left side we get 2a-2c-d=(c-a)m^2 So 2a>2c+d

OpenStudy (pawanyadav):

Yes , sorry ... I'm just confused in my question... You are right

OpenStudy (samigupta8):

Okk..np

OpenStudy (pawanyadav):

I'll tag you in my question .if you can help

OpenStudy (samigupta8):

I replied lol..